OpenStudy (anonymous):

Determine the quadratic equation for the parabola: minimum value: -5; x-intercepts: -2 and 3

OpenStudy (anonymous):

y = a(x+2)(x-3)

OpenStudy (anonymous):

get dy/dx = 0 set the minimum value =-5 and find a

OpenStudy (amistre64):

i might be off on that lol

myininaya (myininaya):

f(x)=a(x-h)^2+k vertex is (n,-5) x-intercepts (-2,0) and (3,0) f(-2)=0 f(3)=0 f(h)=-5=k f(-2)=a(-2-h)^2+k=0 f(3)=a(3-h)^2+k=0 a(4+2*2h+h^2)+k=0 a(h^2+4h+4)+k=0 => a(h^2+4h+4)=-k f(3)=a(3-h)^2+k=0 a(9-2*3h+h^2)+k=0 a(9-6h+h^2)+k=0 => a(9-6h+h^2)=-k so a(h^2+4h+4)=a(9-6h+h^2) h^2+4h+4=h^2-6h+9 4h+6h=9-4 10h=5 h=2 h=2 now we need to know a? a(3-h)^2+k=0 a(3-2)^2-5=0 a-5=0 a=5 so f(x)=5(x-2)^2+5

myininaya (myininaya):

f(x)=5(x-2)^2-5*

myininaya (myininaya):

wait my (-2,0) is not working :(

OpenStudy (amistre64):

get the generic family by (x+2)(x-3); the point in the middle of your zeros is the vertex :) so use that point to calibrate this to -5

OpenStudy (amistre64):

-2+3 = 1; /2 = 1/2 the point (1/2,y) should equal be (1/2,-5) when youre done

myininaya (myininaya):

omg i see my mistake

myininaya (myininaya):

im such a tard 10h=5 h is not 2 h is 1/2

OpenStudy (amistre64):

(1/2)^2 -(1/2) -6 = at x=.5; we get -6.25; -5/-6.25 = the scalar perhaps?

OpenStudy (amistre64):

I get .8 if its right; now multiply everything in the equation by .8 to see if it fits :)

myininaya (myininaya):

a(3-1/2)^2-5=0 a(3-1/2)^2=5 a(5/2)^2=5 a(25/4)=5 a=4/25*5 a=20/25 a=4/5 so f(x)=4/5*(x-1/2)^2-5

OpenStudy (amistre64):

.8 x^2 -.8 x -4.8 is what I get :)

myininaya (myininaya):

i win :) my works does yours

myininaya (myininaya):

your x^2 coefficent is right...

myininaya (myininaya):

looks good amistre :)

OpenStudy (amistre64):

mine works lol; I just tested it :)

myininaya (myininaya):

i think i approached this from 3 blocks away when i could had approached it if i just stayed in my own house

myininaya (myininaya):

you like my metaphor

myininaya (myininaya):

?

OpenStudy (amistre64):

it ok; needs a rocket tho lol

myininaya (myininaya):

lol

myininaya (myininaya):

later